a rectangular football field is 20 feet longer than twice the width. perimeter of the field is 160 feet. find the dimension of the field.
Found 3 solutions by Boreal, MathTherapy, ikleyn:
Answer by Boreal(15235)

width is x
length is 2x+20
P=2 length+ 2 widths
160=2(x-20)+2x
80=x+20+x
2x+20=80
x=30 feet width
x+20=50 feet length

Answer by MathTherapy(10551)


a rectangular football field is 20 feet longer than twice the width. perimeter of the field is 160 feet. find the dimension of the field.
Correct answer: 
IGNORE any-one who says otherwise!

Answer by ikleyn(52778)

.
width is x

length is 2x+20

P = 2*length+ 2*widths

160 = 2(2x+20) + 2x

80 =  2x+20 + x


x =  20 feet width

2x+20 = 60 feet length